Multi-Lever company producers Glayde, a popular room deodorizer. Unfortunately, the production process releases sulfur dioxide into the atmosphere. The marginal private cost (MC) of producing Glayde is MC=Qs, The marginal social cost (MSC) is MSC=1.5Qs, The demand curve for Glayde (there are no external benefits) is P=12-0.5Qd. 7.1.In an unregulated market, what is the equilibrium quantity of Glayde? 7.2. To achieve allocative efficiency, what should be the equilibrium quantity of Glayde? 7.3. Compare and describe the unregulated market quantity and price with the allocatively efficient quantity and price. 7.4. The government wants to impose a tax on Glayde to achieve the allocatively efficient quantity of output. What should the tax be?
Answer #1

(7.1)

In unregulated equilibrium, P = MC.

12 - 0.5Q = Q

1.5Q = 12

Q = 8

P = Q = 8

(7.2)

In efficient equilibrium, P = MSC.

12 - 0.5Q = 1.5Q

2Q = 12

Q = 6

P = 1.5 x 6 = 9

(7.3)

Compared to allocative efficient outcome, in unregulated equilibrium,

Quantity is higher by (8 - 6) = 2

Price is lower by (9 - 8) = 1.

(7.4)

When Q = 6,

MC = Q = 6

MSC = 1.5 x 6 = 9

Unit tax = MSC - MC = 9 - 6 = 3 per unit